When the price of a product rises from RS 100 to RS 200, quantity supplied increases from 50 units to 80 units. What is the value of price elasticity of supply?
1
Expert's answer
2020-09-28T08:02:46-0400

The value of price elasticity of supply is:

"Es = \frac{80 - 50} {200 - 100} × \frac {200 + 100} {80 + 50} = 900/1300 = 0.69."

So, the supply is inelastic.